OpenStudy (anonymous):

4.Compare P (AnB) with P (A U B), and explain what each probability means in the context of the situation and data you collected. Table: boys girls homework 58 62 120 no homework 23 22 45 81 84 165

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

define event A and event B

OpenStudy (anonymous):

A=No homework and B=Girls

OpenStudy (amistre64):

Are A and B independant or dependant?

OpenStudy (anonymous):

this is what i did for question 3. P(A l B) = 22/84 ≈0.26=26% 26% of no homework represent the girls. P(A l B)=22/45=0.04= 4% 4% of the no homework is girls. No we can see that P (A l B) ≠ P (B l A).

OpenStudy (amistre64):

P(homework | girls) is not 22 out of 84, its 62 out of 84

OpenStudy (amistre64):

P(homework|all events) = 120 out of 165

OpenStudy (anonymous):

i was doing the no homework number

OpenStudy (anonymous):

*numbers

OpenStudy (amistre64):

... these old eyes of mine lol

OpenStudy (anonymous):

lol, its okay :)

OpenStudy (amistre64):

fine ... P(no hw | girls) = 22/84 P(no hw | all events) = 45/165 since the P(no hw) changes depending on the given condition; P(no hw) is not independant

OpenStudy (anonymous):

For question 3 the second answer equation is B l A not A l B

OpenStudy (anonymous):

oaky so it would be 3%?

OpenStudy (amistre64):

\[P(A|B)=\frac{P(AnB)}{P(B)}\] \[P(A|B)~P(B)={P(AnB)}\] if P(A) is independant of a given condition, then P(A|B) = P(A) \[P(A)~P(B)={P(AnB)}~\text{, if}~P(A|B)=P(A)\]

OpenStudy (anonymous):

could u show me with the numbers in place?

OpenStudy (amistre64):

P(AuB) = P(A) + P(B) - P(AnB) so the determing the independant nature if useful in some aspects P(AuB) = P(A) + P(B) - P(A|B) P(B)

OpenStudy (amistre64):

P(no hw | girls) = 22/84 P(no hw | all events) = 45/165 the P(no hw) is not independant P(no hw and girls) = 22 out of 165 or stated another way P(no hw and girls) = P(no hw| girl) * P(girls) = 22/84 * 85/165

OpenStudy (amistre64):

84/165 that is

OpenStudy (anonymous):

im so sorry, open study is acting up...

OpenStudy (anonymous):

or maybe its just my computer

OpenStudy (anonymous):

so it would come out as 4%

OpenStudy (amistre64):

P(AnB) = 22/165 = .1333, or 13.33%

OpenStudy (anonymous):

yea. i was just finna write that. didnt do the whole problem

OpenStudy (amistre64):

P(AuB) = (84+45-22)/165

OpenStudy (anonymous):

its 65%

OpenStudy (anonymous):

u there?

OpenStudy (amistre64):

lets make sure i understand what your question is to start with. can you post it?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

Pose a question for which a two-way table is appropriate. Boys Girls Total Homework 58 62 120 No homework 23 22 45 Total 81 84 165 3. Choose a row and column and compare P (A | B) with P (B | A). Explain what each probability means in the context of the situation and data you collected. Let A=No homework and B=Girls P(A l B) = 22/84 ≈0.26=26% 26% of no homework represent the girls. P(B l A)=22/165=0.27= 3% 3% of the no homework is girls. No we can see that P (A l B) ≠ P (B l A). 4. Compare P (A∩B) with P (A∪B), and explain what each probability means in the context of the situation and data you collected.

OpenStudy (amistre64):

P(B|A) is not right, 22 girls out of 45 no hw, 22/45

OpenStudy (anonymous):

0.48888888888=5%

OpenStudy (amistre64):

no, you are confusing probability with percentage. move 2 decimal places

OpenStudy (amistre64):

.4888 -> 48.88%

OpenStudy (anonymous):

ooo,im really horrible at math.... sorry

OpenStudy (amistre64):

i seem to have worked out 4 in my previous postings :)

OpenStudy (anonymous):

wow!!! u did.... im sorry for wasting ur time!!!!!! thanks for all ur help though, and putting up with me :) @amistre64

OpenStudy (amistre64):

'sok, good luck :)